OpenStudy (anonymous):

Translate the word phrase into a math expression. three times the sum of a number and four. A. 3n + 4 B. 3 + 4n C. 3 + (n + 4) D. 3(n + 4)

OpenStudy (readergirl12):

Its A ^.^

OpenStudy (anonymous):

D

OpenStudy (anonymous):

WAIT

OpenStudy (anonymous):

Which property justifies this statement? If x + y = z and z = x + 4, then x + y = x + 4. A. distributive B. associative C. symmetric D. transitive

OpenStudy (readergirl12):

I think it's B I am not 100% sure though

OpenStudy (anonymous):

CAN U CHECK

OpenStudy (readergirl12):

Yeah its correct

OpenStudy (anonymous):

WAIT

OpenStudy (anonymous):

Translate the word phrase into a math expression. three times the sum of a number and four. A. 3n + 4 B. 3 + 4n C. 3 + (n + 4) D. 3(n + 4)

OpenStudy (readergirl12):

A

OpenStudy (anonymous):

D